Are you likely to purchase an item promoted by a celebrity on a social media site? According to a survey, 26% of social media users have made such a purchase Complete parts (a) through (d) below a Suppose that the survey had a sample size of n = 1000 Construct a 95% confidence n er a estima purchased an item promoted by a celebrity on a social media site. or the po a o opo on of soc al media use s (Type an integer or a decimal. Round to four decimal places as needed.)19A telecommunications company wants to estimate the proportion of households that would purchase an additional telephone line if it were made available at a substantially reduced installation cost. Data are collected from a random sample of 500 households. The results indicate that 155 of the households would purchase the additional telephone line at a reduced installation cost. Complete parts (a) and (b) below. a. Construct a 95% confidence interval estimate for the population proportion of households that would purchase the additional telephone line (Round to four decimal places as needed.)
Answer #1

Solution :

n = 500

x = 155

\hat p = x / n = 155 / 500 = 0.31

1 -\hat p = 1 - 0.31 = 0.69

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.960

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.960* (\sqrt((0.310 * 0.690) / 500)

= 0.0405

A 95 % confidence interval for population proportion p is ,

\hat p - E < P <\hat p + E

0.310 - 0.0405 < p < 0.310 + 0.0405

0.2695 < p < 0.3505

The 95% confidence interval for the population proportion p is : ( 0.2695 , 0.3505)